# Amplifying Main Memory-Based Timing Covert and Side Channels using Processing-in-Memory Operations

### v2

@@ 5.2 Validation of Proof-of-Concept
We first demonstrate the PoC results and how we successfully mount the attack in the simulated system. Figures 13 and 14 show the latency (in cycles) measured by the receiver after executing a PiM instruction, one for each of the 16 banks (we assume that the message is 16-bits large), in IMPACT-PnM and IMPACT-PuM, respectively. As we observe, the receiver is able to successfully distinguish whether the sender causes contention in the bank in both IMPACT-PnM and IMPACT-PuM. As shown in the figure, given a latency threshold of 150 cycles, the receiver can decode the complete message: if the latency is less than 150 cycles (i.e., row buffer hit), the receiver decodes the message as a logic-0, otherwise (i.e., row buffer conflict) it decodes the message as a logic-1.
Figure 13: Measured latency (in cycles) by the receiver to execute a PEI instruction for each one of the 16 banks in order to decode a 16-bit message.
@@
Figure 14: Measured latency (in cycles) by the receiver to execute a RowClone instruction for each one of the 16 banks in order to decode a 16-bit message.
5.3 Throughput Analysis of IMPACT Covert-Channel Attacks
@@
We examine the throughput of all different attacks to understand the potential of IMPACT in establishing a high throughput covert channel between the sender and the receiver. Figure 15 shows the throughput achieved by all different attacks when one sender and one receiver thread run concurrently across systems with increasing LLC sizes. We make four key observations. First, IMPACT-PnM and IMPACT-PuM achieve significantly higher throughput than all other attack vectors. For example, IMPACT-PnM achieves a throughput of 12.87 Mb/s irrespective of the cache size, which is 2.29$\times$ (up to 4.91$\times$ with an LLC size of 128MB) higher than the throughput achieved by the second-best performing attack, DRAMA-clflush. Second, DRAMA and Streamline achieve decreasingly lower throughput as the LLC size increases. This is because the larger the LLC size, the higher the latency it takes to access the LLC. Third, IMPACT-PuM achieves 10% higher throughput than IMPACT-PnM since the sender can issue all RowClone requests in parallel, whereas in IMPACT-PnM the sender has to send the PiM instructions in a sequential manner. Fourth, the DMA engine attack achieves a throughput of 5.27 Mb/s irrespective of the cache size, but is almost 2.4x slower than the PnM-based one due to the OS-related overheads associated with it (e.g., context switch overhead, OS instructions, etc.). Fifth, the PnM-OffChip attack leads to high throughput, up to 12.64 Mb/s, but the throughput gets reduced down to 10.64 Mb/s as the cache size increases since the off-chip predictor consults more in favor of caching data when the LLC is large, essentially reducing the attack throughput. We conclude that IMPACT-based attacks can achieve significantly higher throughput than the existing state-of-the-art cache covert channels.
Figure 15: Communication throughput comparison across increasing LLC sizes.
Breakdown of Throughput Improvements. The impact of the fundamental difference between PiM architectures and traditional architectures can be seen as the difference between PnM-OffChip and DRAMA-clflush attack throughputs in Fig. 15. The impact of bypassing the locality monitor can be seen as the difference between IMPACT-PnM and PnM-OffChip attack throughputs. Finally, the impact of the parallelism provided by RowClone can be seen as the difference between IMPACT-PuM and IMPACT-PnM attack throughputs.
@@
To better understand why IMPACT-PuM achieves higher throughput than IMPACT-PnM, we evaluate the time it takes for the sender and the receiver to perform their respective operations in both attacks. Figure 16 shows the time (in cycles) it takes for the sender to send the message and for the receiver to read the message in IMPACT-PnM and IMPACT-PuM. As we observe, in IMPACT-PuM, the sender can send the message 14$\times$ faster than the sender in IMPACT-PuM since it can issue all RowClone requests in parallel. In IMPACT-PnM, even though the sender is significantly slower than the sender of IMPACT-PuM, the sender and the receiver overlap, at the cost of synchronization, in order to hide the latency of the sender. This is why IMPACT-PnM achieves only 10% lower throughput than IMPACT-PuM.
@@
5.4 Throughput Analysis of IMPACT Side-Channel Attack
Figure 17 shows the throughput (taking into account only correct identification of activated hash table entries), left y-axis, and the error rate (i.e., wrong identification of activated hash table entries), right y-axis, of the side-channel attack against DNA-RM across different number of DRAM banks. We make two key observations with respect to the throughput of the attack. First, when using a PiM-enabled DRAM device with 1024 banks, the attacker is able to leak information about the sample genome at a throughput of 7.57 Mb/s. Second, as the number of banks increases, even though the throughput of the attack decreases down to 2.56 Mb/s (since the attacker has to probe more banks and the error rate also increases), the leaked information provides more exact information about the sample genome since the amount of hash table data per bank decreases, reducing the number of potential locations of the reads in the reference genome.
@@ 5.4 Throughput Analysis of IMPACT Side-Channel Attack
We make two key observations with respect to the error rate of the attack. First, for the 1024 bank configuration, the error rate of the attack is less than 5%. Second, we observe that the error rate of the attack is less than 15% across all different DRAM bank configurations. As the number of banks increases, the error rate of the attack increases since the attacker has to probe more banks, which makes the attack more prone to noise and reduces its bandwidth(i.e., DRAM command delay, memory controller scheduling, etc.).
Even though the error rate of the attack increases as the number of banks increases, the attacker is able to leak information about the test genome with higher accuracy. For example, when using a PiM-enabled DRAM device with 1024 banks, the victim activates a row with 16 hash table entries, and the attacker is able to identify if the sample read is present in any of the locations pointed by the 20 hash table entries with an accuracy of 96%. When using a PiM-enabled DRAM device with 2048 banks, the victim activates a row with 8 hash table entries, and the attacker is able to identify if the sample read is present in any of the locations pointed by the 8 hash table entries with an accuracy of 91%. The exact end-to-end accuracy of identifying the properties of the sample genome depends on the chaining mechanism and the imputation algorithm [117, 118] used by the attacker, and it is out of the scope of this work.
@@
Figure 17: Leakage throughput and error rate of the side-channel attack against DNA-RM across different numbers of DRAM banks.
6 Mitigations for IMPACT
This section discusses the effectiveness of three different techniques to mitigate IMPACT: (i) An address mapping scheme that partitions memory banks across different processes, (ii) a memory controller policy that enforces closing the row buffer after every memory access, and (iii) a memory controller policy that enforces constant-time DRAM accesses to hide the timing differences from the attackers.
Memory Partitioning (MPR). Memory bank partitioning can serve as an effective countermeasure against main-memory-based covert channel attacks. By dividing the memory into separate banks, each with its own dedicated resources and access control, memory bank partitioning restricts the interactions between different memory regions. This partitioning prevents unauthorized communication channels from crossing the boundaries of the designated banks. In the context of a IMPACT, memory bank partitioning limits the sender’s ability to transmit information across different memory banks and the receiver’s ability to observe contention in different memory banks. By segregating the memory into distinct banks, the covert channel’s ability to exploit timing differences is significantly constrained. However, statically partitioning DRAM resources across applications disables data sharing, which can significantly (i) increase data duplication, (ii) increase memory underutilization, and (iii) lead to frequent swapping from the disk when the memory space is full.
@@ 6 Mitigations for IMPACT
Closed Row Policy (CRP). Enforcing a memory controller policy that closes a DRAM row after each operation can be an effective measure to mitigate and eventually stop timing covert channel attacks. By closing the DRAM row after each operation, the memory controller ensures that subsequent memory accesses, whether legitimate or covert, cannot access the same row without incurring the overhead of row activation. In the context of IMPACT, which heavily relies on the difference in memory access latency to transmit information, closing the DRAM row after each operation disrupts the timing patterns exploited by the attacker. However, as we show in Section 6, enforcing a closed row policy leads to significant performance overheads, even for workloads with low data locality, as every DRAM access will lead to a DRAM row buffer miss.
@@
Constant-Time DRAM Accesses (CTD). One way to mitigate main-memory-based covert channel attacks is to enforce a constant-time DRAM access policy. When a memory access is issued, the memory controller waits for a fixed amount of time before returning the data to the processor. This fixed amount of time is equal to the worst-case DRAM access latency. By enforcing a constant-time DRAM access policy, the memory controller ensures that all memory accesses take the same amount of time, regardless of the accessed memory location. In the context of IMPACT, where the communication relies on the timing or side effects of memory accesses, enforcing a constant-time DRAM access policy obfuscates the timing patterns exploited by the attacker. However, as we show in Section 6, enforcing a constant-time DRAM access policy leads to significant performance overheads as every DRAM access costs the worst-case DRAM access latency.
@@
Performance Evaluation. To better understand the effectiveness of the countermeasures, we evaluate the performance of CRP and CTD compared to the baseline open-row policy for 5 multi-programmed graph applications which (i) use and share the input graph (i.e., accesses to the same DRAM bank) and (ii) experience a mix of low and high data locality, i.e., the number of last-level cache misses per kilo instructions. We use a 2-core system where each core runs a different instance of the same application (e.g., core 0 runs an instance of PageRank and core 1 the same instance of PageRank using the same input). We evaluate five workloads from GraphBIG benchmark suite [123]: Betweenness Centrality (BC; Data set size: 7.8 GB; LLC MPKI: 0.57), Breadth-First search (BFS; 7.6 GB; 38.59), Connected components (CC; 7.9 GB; 45.2), Triangle counting (TC; 8.2 GB; 5.08), PageRank (PR; 7.0 GB; 1.86).
@@
Figure 18 shows the performance overhead of CRP and CTD compared to the open-row policy for all 5 workloads. We make two key observations. First, CTD leads to the highest performance degration, 26% on average, compared to the open row policy. Especially, in workloads like PR and BC, CTD leads to a performance degradation of 37% and 45% respectively. Second, CRP leads to a performance degradation of 15% on average compared to the open row policy but leads to low performance overheads in workloads that do not benefit from the open row policy like TC, CC and BFS. We conclude that existing countermeasures lead to high performance overheads and there is a dire need to come up with techniques that improve security with low performance overhead.
@@ 7 Related Work
To the best of our knowledge, this is the first work that evaluates the implications of PiM on timing attacks.
@@
Security Implications of PiM Systems. A set of prior works [127, 128, 129] discuss specific security vulnerabilities in PiM systems. Arafin et al. [127] survey and theoretically analyze security vulnerabilities in PiM systems, with a focus on five specific attack scenarios. In this work, we quantitatively and qualitatively analyze and evaluate main memory-based covert and side channels in detail. Wang et al. [129] extracts neural network architectural information from an RRAM accelerator using power side channels. Similarly, Ensan et al. [128] proposes an attack revealing implemented intellectual property in RRAM-based PiM architectures using power/timing side channels. In contrast, in this work, we demonstrate how to establish high-throughput timing covert channels using PiM operations through software.
Timing Attacks. A significant body of work has focused on exploring timing attacks [130, 131]. These attacks involve exploiting variations in the execution time of memory and processor core operations to infer sensitive information. Researchers have investigated various timing attack techniques, such as cache-based attacks [132, 133, 134, 121, 135, 136, 137, 138, 139] to uncover vulnerabilities in traditional systems. There is an extensive body of research related to traditional covert-channel attacks, including Flush+Reload [136], where the sender flushes a shared cache line and monitors the access time by the receiver to infer transmitted data. Flush+Flush [135] utilizes cache flush operations to probe for the presence of specific data patterns. Prime+Probe [140] exploits cache set conflicts to communicate information between sender and receiver. Streamline [121] is a flush-less cache covert channel attack that enables asynchronous communication. We compare IMPACT to Streamline in Section 5.3 and show that IMPACT achieves significantly higher throughput.
@@
Other Main Memory-based Timing Attack Mitigations. We already extensively discuss and evaluate mitigations that fundamentally eliminate the timing channel in Section 6. We also discuss practical solutions that restrict access to the timing channel through cache management methods and show that they are not applicable to PiM-based attacks. Another mitigation method is to restrict the usage of fine-grained timers to prevent attackers from exploiting the timing channel. A set of modern processors limits access to fine-grained timers such as Apple M1 cores [141]. Fine-grained timers are used in a wide range of user space applications, such as performance and power consumption monitoring, multithreading, synchronization and locking, optimizations on protocols such as TCP. Restricting access to fine-grained timers can also disable these applications or degrade their effects.
DAGguise [142] is a defense mechanism against memory timing side-channel attacks. DAGguise utilizes a memory access pattern representation obtained through offline profiling and shapes an application’s memory access pattern to eliminate information leakage. However, DAGguise does not mitigate row buffer-based timing channels as is and requires the memory controller to employ a closed row policy (evaluated in §6).
@@
Enhancing System Security using PiM. A set of prior works proposes leveraging PIM to enhance the security of modern computing systems. SecNDP [143] proposes an encryption and verification mechanism for untrusted PiM devices that enables computation over encrypted data. SecNDP only focuses on protecting confidentiality and does not consider or analyze potential timing channels in contrast to IMPACT. Nelson et al.[144] propose offloading security critical tasks to near-memory compute units to avoid storing sensitive information in caches and eliminate cache-based side-channel attacks leaking them. In contrast, in this work, we demonstrate and evaluate the implications of PiM on main memory-based timing attacks. InvisiMem. [145] and ObfusMem [146] propose leveraging 3D-stacked memory to design secure processors that provide oblivious RAM equivalent guarantees.
@@ 8 Conclusion
The adoption of processing-in-memory (PiM) architectures has gained significant attention due to their ability to provide high performance and low energy consumption by reducing data movement bottlenecks. In this work, we introduce IMPACT, a new high-throughput main memory timing attack that takes advantage of the direct memory access provided by PiM architectures. IMPACT exploits the shared DRAM row buffer to 1) establish covert channels to communicate with other processes at high throughput and 2) leak security-critical information of victim processes based on their memory accesses. IMPACT covert channels leverage parallelism and eliminate cache bypassing steps, resulting in increased throughput compared to existing main memory-based covert channels. IMPACT side-channel attack targets a genomics application (i.e., DNA read mapping) and leaks the private characteristics of a user’s sample genome by leveraging PiM operations. The proposed countermeasures for IMPACT and the discussion on secure and robust PiM architecture designs highlight the need for addressing security concerns in PiM systems.

### v3

@@ 5.2 Validation of Proof-of-Concept
We first demonstrate the PoC results and how we successfully mount the attack in the simulated system. Fig. 7 shows the latency (in cycles) measured by the receiver after executing (a) a PEI instruction and (b) a RowClone operation, one for each of the 16 banks assuming the message is 16-bits large, in IMPACT-PnM and IMPACT-PuM, respectively. We observe that the receiver can successfully distinguish whether the sender causes contention in the bank in both IMPACT-PnM and IMPACT-PuM. The receiver can decode the complete message using a threshold value of 150 cycles as follows. If the latency is less than 150 cycles (i.e., row buffer hit), the receiver decodes a bit as logic-0, otherwise (i.e., row buffer conflict) it decodes the bit as logic-1.
@@
Figure 7: Latency (in cycles) measured by the receiver to execute (a) a PEI instruction and (b) a RowClone instruction for each bank in order to decode a 16-bit message.
5.3 Covert-Channel Attack Throughput Analysis
@@
Fig. 8 shows the throughput achieved by different covert-channel attacks when one sender and one receiver thread run concurrently across systems with increasing LLC sizes. We make five key observations. First, IMPACT-PnM and IMPACT-PuM achieve significantly higher throughput than all other attack vectors. IMPACT-PnM and IMPACT-PuM achieve communication throughputs of 12.87 Mb/s and 14.16 Mb/s, respectively, irrespective of the cache size. They support up to 4.91$\times$ and 5.41$\times$ of the throughput achieved by the state-of-the-art main memory-based attack, DRAMA-clflush. Second, IMPACT-PuM achieves 10% higher throughput than that of IMPACT-PnM since the sender can issue all RowClone requests in parallel, whereas in IMPACT-PnM, the sender has to send the PiM instructions in a sequential manner. Third, as the LLC size increases, the throughput of DRAMA-eviction, DRAMA-clflush, and Streamline decreases significantly. This is because the larger the LLC size, the higher the latency it takes to access the LLC. Fourth, the DMA engine attack achieves a throughput of 5.27 Mb/s irrespective of the cache size but is almost 2.4$\times$ slower than IMPACT-PnM due to its OS-related overheads (e.g., context switch overhead, OS instructions). Fifth, the PnM-OffChip attack achieves a maximum throughput of 12.64 Mb/s. As the LLC size increases, the throughput is reduced to 10.64 Mb/s because the off-chip predictor decides to cache more data when the LLC is large. We conclude that IMPACT achieves significantly higher throughput than the existing state-of-the-art main memory-, cache- and DMA engine-based covert-channel attacks.
Figure 8: Throughput across increasing LLC sizes.
@@
Breakdown of Throughput Improvements. Fig. 9 shows the time (in CPU cycles) it takes for the sender to send the message and for the receiver to read the message in IMPACT-PnM and IMPACT-PuM. We observe that in IMPACT-PuM, the sender can send the message 14$\times$ faster than the sender in IMPACT-PuM since it can issue all RowClone requests in parallel. In IMPACT-PnM, even though the sender is significantly slower than the sender of IMPACT-PuM, the sender and the receiver overlap, at the cost of synchronization, to hide the latency of the sender. This is why IMPACT-PnM achieves only 10% lower throughput than IMPACT-PuM.
@@
5.4 Side-Channel Attack Throughput Analysis
Fig. 10 shows the throughput and the error rate of the side-channel attack against read mapping (RM) across different numbers of DRAM banks in the left and right y-axis, respectively. We measure the throughput of the attack based on the correct guesses and calculate the error rate based on the number of incorrect guesses. The exact end-to-end accuracy of identifying the properties of the sample genome depends on the chaining mechanism and the imputation algorithm [110, 111] used by the attacker, and it is out of the scope of this work.
@@ 5.4 Side-Channel Attack Throughput Analysis
We make two key observations. First, with a PiM-enabled DRAM with 1024 banks, the attacker can leak information about the sample genome at a throughput of 7.57 Mb/s and an error rate of $<5$%. Second, as the number of banks increases to 8196, the throughput of the attack decreases to 2.56 Mb/s and the error rate increases to $<15$%. This is because the attacker has to probe more banks, which makes the attack more prone to noise and reduces its bandwidth.
However, at the same time, the leaked information provides more exact information about the sample genome since the number of hash table entries per bank decreases, reducing the number of potential locations of the reads in the reference genome per bank. For example, in a PiM-enabled DRAM device with 1024 banks, the victim activates a row with 16 hash table entries. Assuming the attacker identifies the activated bank correctly, they determine that one of the 16 hash table entries was targeted by the victim. In contrast, in a PiM-enabled device with 2048 banks, the victim activates a row with 8 hash table entries. When the attacker identifies the activated bank correctly, they determine that one of the 8 hash table entries was targeted by the victim, which leaks more precise information on the sample genome.
@@
Figure 10: Leakage throughput and error rate of the RM side-channel attack across different numbers of DRAM banks.
6 Defenses against IMPACT
This section discusses the effectiveness of three defense mechanisms that eliminates the row buffer-based timing channel.
Memory Partitioning (MPR). Bank-level memory partitioning enforces that each DRAM bank can only be accessed by one process that it is allocated to. In IMPACT, this limits the sender’s ability to transmit information across different DRAM banks and the receiver’s ability to observe contention created by the sender. However, this approach has three major drawbacks. First, it can lead to limiting the number of concurrently running applications due to the fixed number of DRAM banks available. Second, it can cause memory underutilization due to allocating DRAM bank-sized memory to applications. Third, it disables data sharing, which can significantly increase data duplication and memory underutilization, and lead to frequent swapping from the disk when the memory space is full.
@@ 6 Defenses against IMPACT
Closed Row Policy (CRP). Enforcing a memory controller policy that closes a DRAM row after each operation can be an effective defense against IMPACT. By closing the DRAM row after each operation, the memory controller ensures that subsequent memory accesses cannot access the same row without requiring a row activation. Closed row policy disrupts the timing patterns exploited by the attacker in IMPACT. However, it also leads to significant performance overheads, even for workloads with low data locality, as every DRAM access leads to a row buffer miss.
@@
Constant-Time DRAM Accesses (CTD). Another defense mechanism is to enforce a constant-time DRAM access (CTD) policy. When a memory access is issued, the memory controller waits for a fixed amount of time before returning the data to the processor. This fixed amount of time is equal to the worst-case DRAM access latency. By enforcing CTD, the memory controller ensures that all memory accesses take the same amount of time, regardless of the memory location. Thus, obfuscates the timing patterns exploited by the attacker in IMPACT. However, this leads to significant performance overheads as every DRAM access has the worst-case latency.
@@
Performance Evaluation. We evaluate the performance of CRP and CTD compared to the baseline open-row policy for 5 multi-programmed graph applications which (i) share the input graph (i.e., access to the same DRAM bank) and (ii) experience a mix of low and high data locality (i.e., the number of last-level cache misses per kilo instructions). We use a 2-core system where each core runs a different instance of the same application (e.g., core 0 runs an instance of PageRank and core 1 the same instance of PageRank using the same input). We evaluate five workloads from GraphBIG benchmark suite [120]: Betweenness Centrality (BC; Data set size: 7.8 GB; LLC MPKI: 0.57), Breadth-First search (BFS; 7.6 GB; 38.59), Connected components (CC; 7.9 GB; 45.2), Triangle counting (TC; 8.2 GB; 5.08), PageRank (PR; 7.0 GB; 1.86).
@@
Fig. 11 shows the performance overhead of CRP and CTD compared to the open-row policy. We make two key observations. First, CTD leads to the highest performance degradation, 26% on average, compared to the open row policy. Second, CRP leads to an average performance degradation of 15% compared to the open row policy but leads to low performance overheads in workloads that do not benefit from the open row policy such as TC, CC and BFS. We conclude that fundamental defense mechanisms that eliminate the timing channel lead to high performance overheads, and there is a dire need for high-performance defense mechanisms.
@@ 7 Related Work
To the best of our knowledge, this is the first work that evaluates the timing attack vulnerabilities in PiM systems.
@@
Security Implications of PiM Systems. A set of prior works [121, 122, 123] discuss specific security vulnerabilities in PiM systems. Arafin et al. [121] survey and theoretically analyze security vulnerabilities, focusing on five specific attacks. In this work, we quantitatively and qualitatively analyze and evaluate main memory-based covert and side channels in detail. Wang et al. [123] extracts neural network architectural information from an RRAM accelerator using power side channels. Ensan et al. [122] proposes an attack on RRAM-based PiM architectures using power/timing side channels. In contrast, we demonstrate how to establish high-throughput timing covert and side channels using PiM operations.
@@
Other Main Memory-based Timing Attack Defenses. We already extensively discuss defenses that fundamentally eliminate the timing channel (§6) and defenses based on cache management methods that are inapplicable. Another approach is to restrict the fine-grained timers’ usage to prevent attackers from measure memory requests’ latencies, which is applied in some modern processors (e.g., Apple M1 cores [124]). However, these timers are used in many user space applications (e.g., performance and power consumption monitoring, multithreading, and synchronization and locking). Restricting access to these timers can disable these applications or degrade their performance. DAGguise [125] is a defense mechanism against memory timing side-channel attacks that obfuscates an application’s memory access pattern. DAGguise does not mitigate row buffer-based timing channels and requires a closed row policy. Trusted execution environments (TEEs) (e.g., Intel SGX [126, 127]) can isolate applications from the rest of the system. However, prior works already showcase multiple side-channel attacks on TEEs leveraging the main memory [128], and other structures [129, 130, 131, 128, 132, 133, 134, 135, 136, 137].
@@
Enhancing System Security using PiM. Several prior works enhance the security by leveraging PiM. SecNDP [138] proposes an encryption and verification mechanism for untrusted PiM devices. It only focuses on protecting confidentiality and does not consider or analyze potential timing channels. Nelson et al. [139] propose offloading security critical tasks to near-memory compute units to avoid storing sensitive information in caches and eliminate cache-based side-channel attacks. In this work, we demonstrate and evaluate the implications of PiM on main memory-based timing attacks. InvisiMem [140] and ObfusMem [141] propose leveraging 3D-stacked memory to design secure processors that provide oblivious RAM equivalent guarantees.
@@ 8 Conclusion
We introduce IMPACT, a set of high-throughput main memory timing attacks that leverage the direct memory access enabled by PiM architectures. IMPACT exploits the shared DRAM row buffer to 1) establish covert channels to communicate with other processes at high throughput and 2) leak security-critical information of victim processes based on their memory accesses. IMPACT covert channels leverage parallelism and eliminate cache bypassing steps, resulting in increased throughput compared to existing main memory-based covert channels. IMPACT side-channel attack targets a genomics application and leaks the private characteristics of a user’s sample genome by leveraging PiM operations. We evaluate and discuss three fundamental defense mechanisms against IMPACT.